OpenStudy (anonymous):

how do you rearrange 0.20=x^2/75-x into a quadratic formula?

OpenStudy (anonymous):

multyply 75-x in each term

OpenStudy (anonymous):

So would it be 0.20(75-x) = x^2 15-0.20x = x^2 then subtract x^2 from each side and get -x^2-0.20x+15 does that sound right?

OpenStudy (anonymous):

RIGHT ,, ALSO YOU CAN DEAL WITH THE EQUATION AS x^2+0.20-15

OpenStudy (anonymous):

thank you